| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  33  
 | 
Thu Mar 24 16:59:35 2022 | 
Rodolfo | Laser | Photodiode signal SO | Fast Photodiode signal at the "SO-Bogen" (South-east 
arc).
 
 | 
     | 
| 
  32  
 | 
Thu Mar 24 13:27:47 2022 | 
Rodolfo | DAQ | Disk space - LXG1297 | In our Linux computer (LXG1297: Laser spectroscopy) 
we have a Drive with 1.8 TB total disk space. 
300 GB are at the moment in use, but I can | 
    | 
| 
  31  
 | 
Wed Mar 23 16:58:33 2022 | 
V. Hannen | Detectors | Status copper mirror and XUV detectors | Over the past days, the parabolic copper mirror 
and XUV detectors haven been prepared and 
tested for the upcoming beamtime (see first | 
      | 
| 
  30  
 | 
Wed Mar 23 16:16:42 2022 | 
V. Hannen | Detectors | Control software for XUV and copper mirror detector | Attached you find a manual for the software 
used to control the movement of the XUV detector 
and the copper mirror detector. 
 | 
    | 
| 
  29  
 | 
Wed Mar 23 15:38:15 2022 | 
Carsten | General | Not all lights in the ESR can be switched off | Three of the lights in the ESR cannot be switched 
off. It seems that these lights are now "emergency 
lights" (with battery?). We will need to | 
   | 
| 
  28  
 | 
Wed Mar 23 11:20:55 2022 | 
Simon | Detectors | PMT operating voltages | During the testing of the PMTs the following 
operating voltages were applied:
 
 | 
   | 
| 
  27  
 | 
Mon Mar 21 21:48:46 2022 | 
Jan | DAQ | MesyTec Shaper Control | The MesyTec Shaper & Constant Fraction (MSCF) 
is used to process the PMT/det signals for 
the ADC in the second DAQ.
 | 
   | 
| 
  26  
 | 
Mon Mar 21 17:03:47 2022 | 
  | DAQ | Changing code of go4 | Correction: Comments in C/C++ start with // 
, i.e. forward slash
 
 | 
   | 
| 
  25  
 | 
Mon Mar 21 16:33:23 2022 | 
  | DAQ | Changing code of go4 | We agreed to indicate changes in the go4 analysis 
code according to the
 
following scheme
 | 
   | 
| 
  24  
 | 
Mon Mar 21 10:46:18 2022 | 
Rodolfo | DAQ | TDC & Scaler description channels | For the First Vuprom we have the following 
channel configuration (Stand:21.03.2021). 
 | 
    | 
| 
  23  
 | 
Mon Mar 21 09:18:28 2022 | 
Rodolfo | Accelerator | Pressure at North-East Arc  | Our vacuum expert has reply. The "IZ Pump" 
in this region was off.
 
 | 
   | 
| 
  22  
 | 
Fri Mar 18 16:58:26 2022 | 
Rodolfo | Accelerator | Pressure at North-East Arc  | The pressure at the north-east arc of ESR, 
where the laser window is located, looks 
suspicious. The sensor E01LF02 shows a value | 
     | 
| 
  21  
 | 
Fri Mar 18 16:51:27 2022 | 
Carsten | DAQ | Main DAQ /MBS error | MBS error "TDC data ready wait" solved. Common 
stop trigger (20 kHz) was missing. (according 
to Sebastian) | 
   | 
| 
  20  
 | 
Fri Mar 18 09:18:38 2022 | 
Rodolfo | Accelerator | Vacuum leak, ODR | After the beaking out process we got a vacuum 
leak in one of the MgF2 windows (Referenz: 
PMT South) of the optical detection region | 
    | 
| 
  19  
 | 
Thu Mar 17 17:56:33 2022 | 
Kristian | Calibration | HV dividers | PTB HV dividers arrived. Divider ratios are:
 
 
PT120:          12003,341 
 | 
    | 
| 
  18  
 | 
Thu Mar 17 17:37:49 2022 | 
Rodolfo | Accelerator | High voltage dividers for E-Cooler | Two dividers, from PTB-Braunschweig, have 
been deliver today. We will use them to monitor 
the electron-cooler voltage in situ.
 | 
    | 
| 
  17  
 | 
Wed Mar 16 12:04:40 2022 | 
Kristian | Detectors | HV cable to particle detector South | Checked HV cable to particle detector south. 
Label is HV2 Scraper. 
 
 | 
   | 
| 
  16  
 | 
Wed Mar 16 08:31:52 2022 | 
Kristian | DAQ | Check cables and timing | Konstantin and I generated a signal in the 
Messhütte and sent it to the ESR. 
 
There, we used a 44m BNC cable that was sufficently | 
   | 
| 
  15  
 | 
Thu Mar 10 15:20:26 2022 | 
Jan  | DAQ | analog branch signal patches to Messhütte  | The additional signals from PMTs and XUV detector 
needed for spectroscopic analysis in the 
second DAQ are patched to AP Messhütte (MH) | 
        | 
| 
  14  
 | 
Thu Mar  3 21:39:21 2022 | 
Rodolfo | Laser | Dye solution, DCM in Ethanol and DMSO | I have prepared a new dye solution of DCM 
in Ethanol and DMSO (1:1). The concentration 
of the "oscillator + preamplifier" is 0,3 | 
    |